### v1

# Weylian.bib

Abstract
Second-order (maximally) conformally superintegrable Hamiltonian systems of abundant type are re-examined, revealing their underlying Weylian geometry. This allows one to naturally extend the concept of c-superintegrability from the realm of conformal geometries to that of Weylian manifolds. The resulting structures, in any dimension, are characterised by conformally invariant equations.
MSC2020: 53C18; 37J35, 70G45.
Keywords: Weylian manifold, second-order superintegrable system.
1 Introduction
Let $M$ be a simply connected, orientable smooth manifold. A Riemannian metric $g$ on $M$ is called a metric. Its associated conformal metric on $M$ is
$c=\left\{\Omega^{2}g~{}|~{}\Omega\in\mathcal{C}^{\infty}(M),\Omega\neq 0\right\}\,.$
We then call $(M,c)$ a conformal manifold. We say that it is flat, if there exists $h\in c$ such that $h$ is flat, in which case all elements of $c$ are conformally flat. The cotangent space $T^{*}M$ of $M$ naturally carries a symplectic structure $\omega=-d\theta$ thanks to the tautological $1$-form $\theta$. It allows us to define the natural vector field $X_{F}\in\mathfrak{X}(T^{*}M)$ associated to a function $F:T^{*}M\to\mathbb{R}$ via
$\omega(X_{F},-)=dF\,.$
Furthermore, this allows one to define a natural Poisson structure on $T^{*}M$, with the Poisson bracket {,}$\{-,-\}:\mathcal{C}^{\infty}(T^{*}M)\times\mathcal{C}^{\infty}(T^{*}M)\to\mathcal{C}^{\infty}(T^{*}M)$ defined by
$\{F_{1},F_{2}\}=\omega(X_{F_{1}},X_{F_{2}})\,.$
In canonical Darboux coordinates $(x,p)$, it takes the form
$\{F_{1},F_{2}\}=\sum_{i=1}^{n}\left(\frac{\partial F_{1}}{\partial p_{i}}\frac{\partial F_{2}}{\partial x^{i}}-\frac{\partial F_{1}}{\partial x^{i}}\frac{\partial F_{2}}{\partial p_{i}}\right)\,.$
Now let $V\in\mathcal{C}^{\infty}(M)$. We call the naturally defined function $H:T^{*}M\to\mathbb{R}$,
$H(x,p)=g^{-1}_{x}(p,p)+V(x)$
a Hamiltonian on $M$. For the purposes of the present paper we also introduce multi-Hamiltonians: given a vector space $\mathcal{V}$ (over $\mathbb{R}$) in the space of functions on $M$, we say that
=$\mathcal{H}=\{g^{-1}_{x}(p,p)+V(x)\in\mathcal{C}^{\infty}(T^{*}M)~{}|~{}V\in\mathcal{V}\}$
is the naturally associated multi-Hamiltonian. We call two multi-Hamiltonians $\mathcal{H}$ and $\mathcal{H}^{\prime}$ conformally related, if there is a function $\Omega\in\mathcal{C}^{\infty}(M)$ such that
=$\mathcal{H}^{\prime}=\{\Omega^{-2}H~{}|~{}H\in\mathcal{H}\}\,.$
We say that a multi-Hamiltonian is non-degenerate if $\dim(\mathcal{V})=n+2$, c.f. [KMK2007, KKM-1, KKM-2, KKM-3] for instance.
Non-degenerate multi-Hamiltonians that are superintegrable, i.e. that admit a certain, large number of independent functions $T^{*}M\to\mathbb{R}$ that Poisson commute with the (full) multi-Hamiltonian, have received considerable attention in the literature, e.g. [KKM2018, Kress&Schoebel, Kress07, Capel_phdthesis, KSV2023, PKM2013]. More precisely, consider non-degenerate, second-order superintegrable systems, which are introduced properly below. These systems are classified in dimensions two and three [Evans1990, KKPM2001, KKM-1, KKM-2, Capel&Kress, Kress&Schoebel]. Cross-relations have been found to so-called quadratic algebras, and İnönü-Wigner algebra contractions have been used to construct examples, e.g. [KM2014, Capel&Kress&Post]. Non-degenerate second-order superintegrability has also been related to hypergeometric polynomials organised in the Askey-Wilson scheme [PKM2013, PKM2011].
A geometric characterisation of these systems has recently been put forward [KSV2023, KSV2024, KSV2024_bauhaus], proposing a tensorial formalism that makes higher dimensions accessible to an efficient investigation. It revealed that the Riemannian metrics underlying such superintegrable systems are often Hessian metrics [AV2024] and that examples on constant curvature spaces arise from solutions to the Witten-Dijkgraaf-Verlinde-Verlinde (WDVV) equation [KSV2024]. Stäckel transformations, also known as coupling constant metamorphosis, are a well-known transformation of second-order superintegrable systems, which naturally embeds into the tensorial framework mentioned earlier [KSV2024, KSV2024_bauhaus]. The first of these references argues that these systems should be considered as structures on conformal manifolds (called c-superintegrability therein). The purpose of the present paper is to clarify this concept by shedding light onto it from the angle of Weylian geometry. As a result, we extend the definition of abundant superintegrable systems to Weylian structures, including the case of $2$-dimensional ones (while the ideas of [KSV2024] are general, there is a focus on $n\geq 3$ regarding the abundant structural equations and c-superintegrability). As we are going to see, c.f. Definitions 4 and 5, subtleties arise if the underlying Weylian geometry is fixed in advance. In contrast, the abundant structure itself defines a preferred Weylian structure, which may differ from a given one. We characterise this difference by invariant data. In Section 2.1, we review and improve the invariant structural equations for dimensions $n\geq 3$, c.f. [KSV2024]. In Section 2.1, we then conclude the paper by determining the analogous invariant equations for the structural equations in dimension $n=2$, based on the structural equations given in [KSV2024_bauhaus], which differ significantly from those in higher dimensions.
1.1 Weylian manifolds
The purpose of the present section is to introduce Weylian manifolds, also known as Weyl manifolds, e.g. [Folland1970, Norden1976, CP1999, BFM2023]. Following [MS2020], we use the attribute Weylian in order to avoid confusion with the concept of Weyl manifolds in Cartan geometry, c.f. [CM2023, CS2009]. The following twofold definition is based on the one given in [MS2020], see also the references therein.
Definition 1.
  1. (i)
    A Weyl structure is given by a triple $(M,c,D)$ consisting of a differentiable manifold $M$, a conformal class $c$ of Riemannian metrics on $M$, and a torsion-free affine connection $D$ satisfying the condition, for $g\in c$,
    $Dg=-\theta_{g}\otimes g$ (1)
    for some differential 1-form $\theta_{g}$ depending on $g$.
  2. (ii)
    A Weylian structure $(M,\Phi)$ is a differentiable manifold $M$ together with a Weylian metric $\Phi$ on $M$, i.e. an equivalence class of pairs $(g,\varphi)$ consisting of a (pseudo-)Riemannian metric $g$ and a real-valued differential $1$-form $\varphi$, identified under
    $(g,\varphi)\sim(\tilde{g},\tilde{\varphi})\quad:\Leftrightarrow\quad\tilde{g}=\Omega^{2}g\,,\quad\tilde{\varphi}=\varphi-d\ln|\Omega|\,,$
    for some function $\Omega\in\mathcal{C}^{\infty}(M)$.
Remark 1.
The two definitions are equivalent as a Weylian metric $\Phi$ gives rise, for each representative $(g,\varphi)\in\Phi$, to a connection $\nabla^{(g,\varphi)}$ satisfying
$\nabla^{(g,\varphi)}g+2\varphi g=0\,.$
This statement goes back to Weyl [Weyl1918, WeylGes]. For more details on these definitions and their equivalence, see [MS2020]. In the remainder of the paper, we are going to speak simply of a Weylian manifold when we consider Weyl(ian) structures and do not intend to specify a particular definition.
We now introduce some basic properties of Weylian manifolds.
Flatness of a Weylian manifold: we say that a Weyl structure is flat, if the underlying conformal metric $c$ is flat. Equivalently, a Weylian structure is flat, if there is $(g,\varphi)\in\Phi$ such that $g$ is flat.
Exactness and closedness of a Weylian manifold: A Weyl structure is exact, if the connection $D$ is the Levi-Civita connection of some metric in $c$, and closed, if this property holds in the neighbourhood of each point. Equivalently, these properties hold, if $\theta_{g}$ is exact or closed, respectively, for some (and then any) element of $c$, c.f. [MP2024, BFM2023]. We define exactness and closedness for Weylian structures accordingly.
1.2 Abundant superintegrable systems
A superintegrable system, or, more precisely speaking, a second-order (maximally) conformally superintegrable system, is a Riemannian manifold $(M,g)$ with a natural Hamiltonian $H:T^{*}M\to\mathbb{R}$ admitting $22n-$ 2 additional functions $F^{(\alpha)}:T^{*}M\to\mathbb{R}$, $21\leq\alpha\leq 2n-$ 2, such that (with $F^{(0)}:=H$):
  1. (i)
    for each $21\leq\alpha\leq 2n-$ 2 there are functions $K^{ij}$ and $W$ on $M$ such that
    $F^{(\alpha)}(x,p)=\sum_{i,j=1}^{n}K^{ij}(x)p_{i}p_{j}+W(x)\,.$ (2)
  2. (ii)
    the functions $F^{(\alpha)}$, $21\leq\alpha\leq 2n-$ 2, satisfy
    $\{F^{(\alpha)},H\}=2\varrho^{(\alpha)}\,H$ (3)
    for suitable $\varrho^{(\alpha)}:T^{*}M\to\mathbb{R}$, respectively. We require these $\varrho^{(\alpha)}$ to be polynomial in the momenta.
  3. (iii)
    the functions $(F^{(\alpha)})_{0\leq\alpha\leq 2n-2}$ are functionally independent.
The first point in the above list is the second-order property, while the number $22n-$ 2 indicates the maximality of the superintegrable system. The condition (3) implies that the functions $F^{(\alpha)}$, $20\leq\alpha\leq 2n-$ 2, are constant when restricted to Hamiltonian trajectories on the zero locus of $H$, i.e. the solution curves $\gamma$ of Hamilton’s system of equations,
$\dot{\gamma}=X_{H}\circ\gamma\,,$ (4)
that in addition satisfy $H(\gamma(t))=0$.
Remark 2.
Note that the condition (3) is a cubic polynomial in the fibre coordinates $p$ (called momenta) on $T^{*}M$. It is easily checked that it has two homogeneous components (in the momenta), namely a cubic and a linear one. A straightforward computation shows that the cubic homogenous part of the polynomial is equivalent to the condition of a conformal Killing tensor,
$\nabla_{X}K(X,X)=\rho(X)g(X,X)\qquad\qquad\forall X\in\mathfrak{X}(M),$
where $K=\sum_{i,j=1}^{n}(\sum_{a,b=1}^{n}g_{ia}K^{ab}g_{bj})\,dx^{i}\odot dx^{k}$, and where
$\rho=\frac{1}{n+2}\left(2\mathrm{div}_{g}(K)+d\mathrm{tr}_{g}(K)\right).$
The linear homogeneous component, on the other hand, can be rewritten as
$dW=\hat{K}(dV)-V\,\rho-dV\wedge\rho\,,$
where $\hat{K}$ denotes the endomorphism naturally associated to $K$ (via the inverse of the metric $g$), and where $\wedge$ is the usual wedge product.
Convention 1.
Without loss of generality, we assume from now on that $K$ (for each $F^{(\alpha)}$, $21\leq\alpha\leq 2n-$ 2, be trace-free. This is no restriction: modifying the trace of a conformal Killing tensor, we again obtain a conformal Killing tensor.
Convention 2.
We will tacitly use a circumflex diacritic to denote, for a symmetric tensor field $Q\in\Gamma(\mathrm{Sym}^{k+1}(T^{*}M))$, the (via $g$) naturally associated section of $\mathrm{Sym}^{k}(T^{*}M)\otimes TM$, whenever the underlying metric $g$ is clear.
We denote by $\Pi_{(ijk)\circ}$ the projector onto the totally symmetric (in $i,j,k$) and trace-free (w.r.t. contraction in any pair of $i,j,k$) part of a tensor field (which may have further indices over which no symmetrisation nor trace-freeness is enforced). The circle $\circ$, in particular, will mean trace-freeness. Consequently, the projector onto the totally tracefree part is denoted by $\Pi_{\circ}$. By $\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$ we mean the space of symmetric and tracefree tensor sections of rank two.
1.3 Abundant systems
In the present paper, we also need the concept of abundant systems. For the following definition, we introduce the $1$-form $\beta\in\Omega^{1}(M)$,
$\beta_{k}$ $=S_{kab}\Xi^{ab}\,.$
Moreover, we denote the norm (with respect to the metric $g$) of a tensor with components $T_{{i_{1}}\cdots{i_{r}}}$ by
$|T|_{g}^{2}=T^{{i_{1}}\cdots{i_{r}}}T_{{i_{1}}\cdots{i_{r}}}\,.$
We may omit the subscript $g$, if the underlying metric is clear.
Definition 2 ([KSV2024]).
Let $(M,g)$ be a conformally flat Riemannian manifold of dimension $n\geq 2$, endowed with a totally symmetric and trace-free tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ and a function $t\in\mathcal{C}^{\infty}(M)$ such that [^1] Compared to [KSV2024_bauhaus], two typographical errors have been corrected in (5fb) and (5fd), involving the replacement of two erroneous signs.
  • if the dimension is $n\geq 3$:
    $\nabla_{i}\bar{t}_{j}$ $=3\,\mathsf{P}_{ij}+\frac{1}{3(n-2)}\left({S_{i}}^{ab}S_{jab}+\frac{n-6}{2(n-1)(n+2)}\,|S|_{g}^{2}\,g\right)$
    $\qquad\qquad\qquad+\frac{1}{3}\left(\nabla_{i}t\nabla_{j}t-\frac{1}{2}|\mathrm{grad}_{g}t|_{g}^{2}\,g\right)$ (5a)
[…]
    00 0 $=\Pi_{\circ}\left(g^{ab}S_{aik}S_{bjl}-g^{ab}S_{ail}S_{bjk}\right)$ (5c)
  • if the dimension is $n=2$:
    $\nabla_{l}S_{ijk}$ $=\Pi_{(ijk)\circ}\left[-\frac{2}{3}S_{ijk}t_{l}+2t_{i}S_{jkl}+\Xi_{ij}g_{kl}\right]$ (5fa)
    $\nabla_{k}\Xi_{ij}$ $=\frac{1}{3}S_{ijk}\,\,|S|_{g}^{2}+\frac{4}{3}\Pi_{(ijk)\circ}\Xi_{jk}t_{i}+\frac{4}{3}\Pi_{(ij)\circ}g_{jk}\beta_{i}$ (5fb)
    $\nabla^{a}\tau_{ak}$ $=-S_{kab}\tau^{ab}+\beta_{k}-\frac{2}{3}\,\Xi_{ka}t^{a}-\frac{4}{9}\,S_{kab}t^{a}t^{b}-\frac{5}{9}\,|S|_{g}^{2}\,t_{k}$
    $\hskip 170.71652pt+\frac{1}{2}\nabla_{k}R-Rt_{k}$ (5fc)
    $\nabla^{2}_{ij}t$ $=\frac{3}{2}\left[\tau_{ij}+\frac{2}{3}S_{ija}t^{a}+\frac{8}{9}\Pi_{(ij)\circ}t_{i}t_{j}-\frac{1}{3}\Xi_{ij}\right]$
    $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{2}g_{ij}\left[\frac{1}{3}\,|S|_{g}^{2}+\frac{3}{2}R\right]$ (5fd)
    We then say that $(S,t)$ is an abundant system on $(M,g)$. If we consider only the underlying smooth manifold $M$ to be fixed, we call $(g,S,t)$ an abundant system on $M$.
    The relationship between abundant systems and superintegrable systems has been elucidated by [KSV2024, KSV2024_bauhaus], and we refer the interested reader to these references for more details: Here we confine ourselves to mentioning that given an abundant manifold, one may integrate a certain system of partial differential equations whose solutions solve (3). In fact, one typically obtains $\frac{1}{2}(n-1)(n+2)$ linearly independent functions $F^{(\alpha)}$ of the form (2) that satisfy (3) for a non-degenerate multi-Hamiltonian. The interested reader will find more details on this construction in [KSV2024, KSV2024_bauhaus] and [KKM2018] as well as in the references therein. In particular, it is shown in [KSV2024] that the equations in Definition 2 are conformally invariant. By this we mean that these equations remain true if we replace $g\mapsto\Omega^{2}g$ for some nowhere vanishing $\Omega\in\mathcal{C}^{\infty}(M)$ (and its Levi-Civita connection accordingly) as well as
    $S\to\Omega^{2}S\,,\qquad t\to t-3\ln|\Omega|\,.$
    Definition 3.
    Two abundant systems $(g,S,t)$ and $(g^{\prime},S^{\prime},t^{\prime})$ on the same smooth manifold are said to be conformally related if $g^{\prime}=\Omega^{2}g$, $S^{\prime}=\Omega^{2}S$ and $t^{\prime}=t-3\ln|\Omega|+c$ for some $c\in\mathbb{R}$.
    2 Abundant systems on Weylian manifolds
    Having introduced abundant systems and Weylian manifolds, we are now going to relate these two structures. Recall that for an exact Weyl structure $(M,c,D)$, the $1$-form $\theta_{g}$ in (1) is exact for any $g\in c$. We thus have $\theta_{g}=d\Theta_{g}$ for some function $\Theta_{g}\in\mathcal{C}^{\infty}(M)$. Analogously, consider the corresponding Weylian structure $(M,\Phi)$. For $(g,\varphi)\in\Phi$, the $1$-form $\varphi$ is also exact, and $\varphi=2d\phi$ for some function $\phi$, c.f. Remark 1. With this terminology, we state the following lemma, which allows us to define abundant systems on Weylian manifolds. Its proof is implied by [KSV2024].
    Lemma 1.
    Let $(M,\Phi)$ be an exact Weylian structure. Let $(g,\varphi)\in\Phi$ and assume that $(M,g)$ admits an abundant system encoded in the tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ and the function $t\in\mathcal{C}^{\infty}(M)$. Next, let $(g^{\prime},\varphi^{\prime})$ be another representative of the Weylian metric $\Phi$.
    Then $(M,g^{\prime})$ carries the abundant system encoded in the structure tensor $S^{\prime}=g^{\prime}g^{-1}S$ and the function $t^{\prime}=t-3(\varphi-\varphi^{\prime})$.
    Note that $g^{\prime}g^{-1}S$ here means that we first raise, then lower one index using the metrics $g$ and $g^{\prime}$, respectively.
    Proof.
    Clearly, $(g^{\prime},\varphi^{\prime})=(\Omega^{2}g,\varphi-\ln|\Omega|)$ for some function $\Omega$. We hence have $S^{\prime}=\Omega^{2}S$ and $t^{\prime}=t-3\ln|\Omega|$. It then follows from [KSV2024] that $(M,g^{\prime},S^{\prime},t^{\prime})$ is an abundant system, given that $(M,g,S,t)$ is one, and since all objects satisfy the properties of conformally related systems. The claim follows. ∎
    We observe that $\hat{S}=g^{-1}S\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and $\hat{S}=(g^{\prime})^{-1}S^{\prime}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ coincide. We also note that in the lemma, $\frac{1}{3}t$ and $\varphi$ transform into $\frac{1}{3}t^{\prime}$ and $\varphi^{\prime}$, respectively, in an analogous manner. The equation $t-3\varphi=t^{\prime}-3\varphi^{\prime}$ hence yields a conformal invariant. We are now ready to define abundant systems on Weylian manifolds.
    Definition 4.
    We call a triple $(M,\Phi,\hat{S})$ consisting of a smooth manifold $M$ with Weylian metric $\Phi$ as well as a (2,1)-tensor field $\hat{S}\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M)\otimes TM)$ a natural abundant Weylian structure if $S=g\hat{S}\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ is trace-free and satisfies the conditions in Definition 2 with $t=3\varphi$.
    We introduce natural abundant Weyl structures and natural abundant Weylian manifolds accordingly, in the obvious way. Note that in the definition, we suppress for the sake of conciseness the irrelevant constant that may be added to $t$ or $\varphi$ (i.e. we identify structures that only differ by such a constant).
    Remark 3.
    Compare Definition 4 to the definition of c-superintegrable systems, c.f. Definition 3.9 in [KSV2024]. We note that our focus here is on the case of abundant systems, but that the underpinning ideas are nontheless of a general nature. Indeed, we observe the fundamental difference that, in the reference, a c-superintegrable system is defined on a merely conformal structure $(M,c)$, i.e. a smooth manifold with a conformal class of metrics. The discussion in the current section has revealed the underpinning Weylian geometry. In [KSV2023], this underlying Weylian structure exists implicitly, c.f. Section 6 of [KSV2024]. For $n\geq 3$, it is indirectly inferred, by Eq. (6.6) of [KSV2023], from a conformal class of abundant superintegrable systems.
    We may reinterpret this as having subsumed the invariant difference function $\mathfrak{t}=t-3\varphi$ bridging between the representative $(g,\varphi)\in\Phi$ of the Weylian metric and the abundant system $(M,g,S,t)$. Indeed, if $\mathfrak{t}\neq 0$, we may replace $\Phi$ be a new Weylian metric $\Phi^{\prime}=[(g,\varphi+\frac{1}{3}\mathfrak{t})]$. In terms of the corresponding Weyl structure this presents itself as a conformal change of the Weylian connection. This observation now allows us to extend abundant systems to pre-fixed Weylian manifolds.
    In the light of the remark, let the Weylian manifold be fixed, e.g. in terms of a Weyl structure $(M,g,D)$. We may then define an abundant system on this Weylian manifold as follows.
    Definition 5.
    Let $(M,\Phi)$ be a Weylian structure. We call $(\hat{S},\mathfrak{t})$ consisting of a (2,1)-tensor field $\hat{S}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and a function $\mathfrak{t}\in\mathcal{C}^{\infty}(M)$ an abundant system on $(M,\Phi)$ if $(M,[(g,\varphi+\frac{1}{3}\mathfrak{t})],\hat{S})$ is an abundant Weylian manifold.
    We hence obtain:
    Theorem 1.
    1. (i)
      An abundant system $(\hat{S},\mathfrak{t})$ on a Weylian manifold $(M,\Phi)$ is a natural abundant Weylian structure precisely if $\mathfrak{t}$ is constant.
    2. (ii)
      If $(\hat{S},\mathfrak{t})$ is an abundant system on the Weylian manifold $(M,\Phi)$, then $(M,[(g,\varphi+\mathfrak{t})],\hat{S})$ is a natural abundant Weylian manifold, $(g,\varphi)\in\Phi$.
    Proof.
    For part (i), observe that the structural equations (5) and (5f), respectively, involve only derivatives of the function $t$.
    2.1 Structural equations for $n\geq 3$
    Our aim is now to give a “representative-free” characterisation of abundant systems on Weylian manifolds. To this end, we need to establish some terminology for a given Weyl structure $(M,g,D)$. In order to do this efficiently, we introduce some terminology. We denote by $\Pi_{\mathrm{Sym}^{3}_{0}}$ the projection onto the trace-free totally symmetric component, and we introduce conformally adapted operators $\mathscr{H}:\mathcal{C}^{\infty}(M)\to\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$,
    $\mathscr{H}(f)=\mathrm{Hess}^{g}(f)-\frac{1}{n}\,g\,\Delta^{g}(f)-\mathring{\mathsf{P}}\,f\,,$
    and the conformal Laplacian
    $\mathscr{L}({f})=-4\,\frac{n-1}{n-2}\Delta^{g}{f}+\mathrm{Scal}^{g}{f}\,.$
    Remark 4.
    To ensure simplicity and accessibility, we apply the usual language of tensor fields and functions on $M$, but mention that an alternative formulation of the structural equations can be put forth in the language of weighted density bundles. The (weight-$w$) density line bundle $L^{w}$ is the bundle whose fibres (at a point $x\in M$) are given by the maps $\ell:\Omega^{n}(M)\to\mathbb{R}$ such that, for $r\in\mathbb{R}\setminus\{0\}$, $\ell(r\alpha)=|r|^{-\frac{w}{n}}\ell(\alpha)$. Similarly, one introduces the weighted tensor bundles $L^{w,k,k^{*}}=L^{w}\otimes(TM)^{k}\otimes(T^{*}M)^{k^{*}}$, for $w,k,k^{*}\in\mathbb{N}_{0}$. These are said to have weight $w+k-k^{*}$[CP1999]. A (tensor) density $\mu$ of weight $\mathsf{w}$ thus transforms as $\mu\to|\Omega|^{\mathsf{w}}\mu$ under conformal rescalings.
    We now introduce the conformally invariant sections $\mathcal{P}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$, $\mathcal{L}\in\mathcal{C}^{\infty}(M)$, for a given Weylian structure $(M,\Phi=[(g,\varphi)])$,
    $\mathcal{P}$ $:=e^{3\varphi}\mathscr{H}\left(e^{-3\varphi}\right)$
    $\mathcal{L}$ $:=e^{3(1-\frac{n}{2})\varphi}\mathscr{L}\left(e^{-3(1-\frac{n}{2})\varphi}\right)\,$
    We furthermore introduce the conformally invariant tensor field
    $\mathscr{S}(X,Y)=\mathrm{tr}(\hat{S}(X,\hat{S}(Y,-)))$
    and its (conformally invariant) trace-free part
    $\mathring{\mathscr{S}}=\mathscr{S}-\frac{1}{n}\,g\,\mathrm{tr}_{g}(\mathscr{S})\,.$
    Next, we introduce
    $\mathfrak{S}(X,Y,Z)=\hat{S}(\hat{S}(Y,Z),X),$
    for $X,Y,Z\in\mathfrak{X}(M)$, and then a tensor field $\mathfrak{A}\in\Gamma(T^{*}M^{\otimes 3}\otimes TM)$ via
[…]
    Note that $\mathfrak{A}$ is conformally invariant. Let $\nabla^{\hat{S}}=\nabla^{g}-\hat{S}$ and denote its curvature and Ricci tensor by $R^{S}$ and $\mathrm{Ric}^{S}$, respectively. We then define the Weyl tensor $\mathrm{Weyl}^{\hat{S}}$ of $\nabla^{\hat{S}}$ (with respect to $g$) by
    $\mathrm{Weyl}^{\hat{S}}(X,Y,Z,W)=g(R^{\hat{S}}(X,Y)Z,W)-(\mathsf{P}^{\hat{S}}\owedge g)(X,Y,Z,W)\,,$
[…]
    $\mathsf{P}^{\hat{S}}=\frac{1}{n-2}\left(\mathrm{Ric}^{\hat{S}}-\frac{\mathrm{tr}_{g}\mathrm{Ric}^{\hat{S}}}{2n(n-1)}\,g\right)\,.$
    We are now ready to characterise abundant Weylian structures.
    Theorem 2.
    Let $(M,\Phi)$ be a flat Weylian structure, $n\geq 3$. Then $(M,\Phi,\hat{S})$, where $\hat{S}$ is a $(1,2)$-tensor field, is abundant, if and only if, for $\alpha\in\Omega^{1}(M)$ and $X\in\mathfrak{X}(M)$,
    $\mathcal{P}$ $=-\frac{1}{9(n-2)}\,\mathring{\mathscr{S}}$ (5g)
    $\mathcal{L}$ $=-\frac{2}{9}\frac{3n+2}{n+2}\,|S|^{2}$ (5h)
    $D_{X}\hat{S}$ $=\mathfrak{A}(X,\cdot,\cdot)$ (5i)
    $\mathrm{Weyl}^{\hat{S}}$ $=0$ (5j)
    Note that (5g)–(5j) are conformally invariant conditions, i.e. that they hold true also after a conformal rescaling.
    Proof.
    The conditions (5g)–(5j) are obviously equivalent to (5). Hence, for each $(g,\phi)\in\Phi$, $(M,g,g\hat{S},t)$, $t=\frac{1}{3}\phi$, defines an abundant manifold. Hence, $(M,\Phi,\hat{S})$ defines an abundant Weylian manifold. ∎
    An analogous statement holds for abundant systems (with non-vanishing $\mathfrak{t}$) on flat Weylian structures, the details of which we leave to the interested reader.
    2.2 Structural equations for $n=2$
    Structural equations can also be written down, in conformally invariant form, for $2$-dimensional abundant Weylian structures. We introduce the tensor section $\mathfrak{Z}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes T^{*}M\otimes TM)$ via
    $\mathfrak{Z}_{ijl}^{k}=g^{ka}\Pi_{(ija)\circ}\left(\Xi_{ij}g_{al}\right)$
    and observe that it is conformally invariant. Moreover, we introduce $\mathfrak{b}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes T^{*}M)$ by
    $\mathfrak{b}_{ijk}=\Pi_{(ij)\circ}\beta_{i}g_{jk}\,,$
    which we likewise observe to be conformally invariant. Reconsidering the relevant equations in Definition (2), we now find the following theorem.
    Theorem 3.
    Let $(M,\Phi)$ be a flat Weylian structure, $n=2$. Then $(M,\Phi,\hat{S})$, where $\hat{S}$ is a $(1,2)$-tensor field, is abundant, if and only if, for $X\in\mathfrak{X}(M)$,
    $D_{X}\hat{S}$ $=\mathfrak{Z}(\cdot,\cdot,X)$
    $D_{X}\Xi$ $=\frac{|S|_{g}^{2}}{3}S(X,\cdot,\cdot)+\frac{4}{3}\mathfrak{b}(\cdot,\cdot,X)$
    $\mathcal{L}^{\mathrm{2D}}$ $=\frac{1}{9}\,|S|^{2}$
    where we introduce
    $\mathcal{L}^{\mathrm{2D}}:=\left(\Delta-\frac{3}{2}R\right)t\,.$
    These conditions are conformally invariant. Note that $\mathcal{L}^{\mathrm{2D}}$ behaves like a density of weight $q=-2$, i.e. $\mathcal{L}^{\mathrm{2D}}\mapsto\Omega^{-2}\mathcal{L}^{\mathrm{2D}}$, under conformal changes $g\mapsto\Omega^{2}g$. Indeed, letting $t^{\prime}=t-\ln|\Omega|$,
    $\Delta t$ $\mapsto\Omega^{-2}\left[\Delta t^{\prime}-2\,dt^{\prime}(\mathrm{grad}_{g}\ln|\Omega|)+2\,g^{-1}(d\ln|\Omega|,dt^{\prime})\right]\,,$
    $R$ $\mapsto\Omega^{-2}\left[R-2\Delta\ln|\Omega|\right]\,,$
    and hence
    $\Delta t-\frac{3}{2}\,R$ $\mapsto\Omega^{-2}\left[\Delta(t-3\ln|\Omega|)-\frac{3}{2}\,(R-2\,\Delta\ln|\Omega|)\right]$
    $=\Omega^{-2}\left[\Delta t-3\Delta\ln|\Omega|-\frac{3}{2}\,R+3\,\Delta\ln|\Omega|\right]=\Omega^{-2}\left[\Delta t-\frac{3}{2}\,R\right]\,.$
    Acknowledgements
    The author thanks Vicente Cortés for discussions. This research was funded by the German Research Foundation (Deutsche Forschungsgemeinschaft) through the Research Grant 540196982. The author also thanks the Forschungsfonds at the University of Hamburg for support.
[…]